OpenStudy (toribyrd33):

If sin Θ = 1 over 4 and tan Θ > 0, what is the value of cos Θ?

OpenStudy (toribyrd33):

negative square root of 15 square root of 15 square root of 15 over 4 negative square root of 15 over 4

Nnesha (nnesha):

make a right triangle  sin = what ? opposite adjacent hypotenuse ?

OpenStudy (toribyrd33):

sin=1/4?

Nnesha (nnesha):

yes that's right but sin cosine and tan \[\rm tan = \frac{ \sin }{ \cos } = \frac{ \rm opposite }{ \rm adjacent }\] \[\rm sin = \frac{ opposite }{ hypothunuse }\] \[\rm cos = \frac{ adjacent }{ hypothunuse }\]

Nnesha (nnesha):